OpenStudy (anonymous):

Solve 4x – 7 – (x + 2) = 3x If the equation is an identity, write “identity”. If there are no solutions, type “none”.

OpenStudy (anonymous):

4x - 7 - x - 2 - 3x = 0 x - x - 7 - 2 = 0 -9 = 0 no solutions

OpenStudy (anonymous):

4x-7-x-2 = 3x 3x-9 = 3x -9 = 3x-3x -9 = 0 so,there is no solutions